### v1

@@ 4.1 Approximate Single-Peaked Matrices
We say a matrix $P$ is $\delta$-approximately single-peaked (or $\delta$-ASP) w.r.t. an order $\prec$ if for every $i,j,l\in[K]$ such that $i\prec j\prec l$ and user $u\in U$, it holds that
$P_{u,j}\geq\min\{P_{u,i},P_{u,l}\}-\delta.$ (2)
@@
The $\delta$-ASP condition proves useful beyond the strict $\delta=0$ case. In fact, it follows directly from (2) that if $P$ is SP w.r.t. $\prec$ and $\tilde{P}$ satisfies $\lVert P-\tilde{P}\rVert_{\infty}\leq\delta$, then $\tilde{P}$ is $2\delta$-ASP w.r.t. $\prec$. This implies that a noisy estimate of an SP matrix remains ASP. Conversely, given any $\delta$-ASP matrix and its ASP order, we can construct a matrix that is (exact) SP with respect to the same order, with each entry differing from the original by at most $\delta$.
@@
We prove Lemma 4 by construction, which can be implemented in $O(UK)$ time. The detailed construction is given in Appendix B. Lemma 4 enables us to work with SP matrices even when the estimate matrix is not SP, while not sacrificing the approximation quality. However, Lemma 4 requires the ASP order $\prec$, which we do not have.
@@ 4.2 Extract-Order Procedure
We start with some intuition. Finding an order boils down to solving a set of contiguity constraints—subsets of arms that must be contiguous in any valid order. To illustrate, fix an SP matrix $P$ and suppose there exists a user $u$ and a partition of $[K]$ into two sets $A_{1}$ and $A_{2}$. Let $\beta=\min_{k\in A_{1}}P_{u,k}$ and $\gamma=\max_{k\in A_{2}}P_{u,k}$. If $\beta-\gamma>0$, then every arm in $A_{1}$ is preferred by user $u$ over every arm in $A_{2}$. In this case, any order $\prec$ that places an arm $j\in A_{2}$ between two arms $i,l\in A_{1}$ violates the 0-ASP condition ((2)). Hence, according to Proposition 3, $\prec$ cannot be an SP order of $P$. Thus, any SP order of $P$ places the arms in $A_{1}$ in a contiguous block. As a result, any order that satisfies all such contiguity constraints is an SP order of $P$.
@@
The Extract-Order algorithm, described in Algorithm 1, formalizes this approach using a PQ tree [6], an efficient data structure that specializes in encoding and resolving contiguity constraints. PQ trees maintain sets of items that must appear consecutively in any valid order, and can be used both to determine whether a consistent order exists and to efficiently construct such an order. Extract-Order initializes a PQ tree (Line 3), and for each user $u$, it sorts the arms by preference values (Lines 4-5). Whenever there is a significant gap (exceeding $2\varepsilon$) between consecutive preference values, which exactly corresponds to the case described in (3), it adds a contiguity constraint requiring all higher-valued arms to appear consecutively (Line 7). Finally, the algorithm returns an order on the columns of $\tilde{P}$ that satisfies all the imposed contiguity constraints, if such an order exists (Line 9). Otherwise, it fails and returns nothing (Line 11)
Notably, the constraints enforced by this procedure are not only consistent with the SP order of $P$, but also ensure bounded valley depth (in the sense of (2)). We formalize this in the following lemma.
@@
Let $\tilde{P}$ be a matrix such that $\|\tilde{P}-P\|_{\infty}\leq\varepsilon$ for some SP matrix $P$. Extract-Order($\tilde{P},\varepsilon$) returns an order $\prec$ such that $\tilde{P}$ is $\left(2K\varepsilon\right)$-ASP w.r.t. $\prec$.
@@ 4.3 Explore-then-Commit Algorithm
4:Construct an SP matrix $\tilde{\Theta}$ from $(\bar{\Theta},\prec)$ via Lemma 4 with $\delta=2K\sqrt{\tfrac{2\ln T}{N}}$
@@
We now present the EMC algorithm, which we implement in Algorithm 2, that combines the tools developed in the previous subsections to achieve sublinear regret. It operates in five phases. First, given an input parameter $N$, the algorithm explores by pulling each arm $k\in[K]$ for $N$ rounds and collects empirical means $\bar{\Theta}$. Second, it applies Extract-Order and extracts an order $\prec$ from $\bar{\Theta}$. The parameter $\varepsilon$ is chosen so that $\lVert\Theta-\bar{\Theta}\rVert_{\infty}\leq\varepsilon$ holds with high probability. In the third phase, it uses $\bar{\Theta}$ and the order $\prec$ to construct the matrix $\tilde{\Theta}$, which is SP w.r.t. $\prec$ and is element-wise close to $\bar{\Theta}$. Fourth, it executes SP-Matching on $\tilde{\Theta}$ to find an optimal matching $\tilde{\pi}$. And in the fifth and last phase, it exploits—it plays $\tilde{\pi}$ for the remaining TNKT-NK rounds. Importantly, the computational effort in Lines 35 is only $O(K^{2}U+K^{2}B)$, ensuring that the algorithm is computationally efficient. This is in sharp contrast to the exponential runtime we obtain for general instances in Section 2.2.
The regret analysis for the commitment phase hinges on the approximation quality between the matrix $\tilde{\Theta}$ used for the matching and the true expected reward matrix $\Theta$. Since $\tilde{\Theta}$ is constructed to be close to the empirical estimates $\bar{\Theta}$, and $\bar{\Theta}$ concentrates around $\Theta$ with sufficient exploration, the value gap $V(\tilde{\pi};\tilde{\Theta})-V(\tilde{\pi};\Theta)$ is bounded by the estimation errors and the approximation errors from Lemma 4 and Extract-Order. By optimizing the choice of exploration rounds $N$, we balance the exploration cost against the quality of the resulting approximation.
@@
Moreover, the computational phases in Lines 3-5 take $O(K^{2}U+K^{2}B)$, which makes it feasible in practical scenarios. This is in sharp contrast to the exponential runtime we obtain for general instances in Section 2.2.
5 Known SP Structure
In this section, we leverage additional structural information to design an efficient algorithm with improved regret guarantees. Specifically, we consider the setting in which both the SP order and user peaks are known. Although in standard stochastic bandit problems knowledge of each user’s peak (i.e., the optimal arm) is sufficient to eliminate regret, here, matching users to their peaks can violate budget constraints. Consequently, identifying the optimal matching requires exploration of the arms’ cardinal utilities.
These structural assumptions enable an optimistic, UCB-style algorithm that achieves a regret bound of $\tilde{O}(U\sqrt{TK})$—a rate that is comparable to our lower bound in Section 6. Our approach involves the construction of optimistic reward estimates using confidence sets [31, Chapter 8.4], as detailed in Section 5.1, while preserving the single-peaked structure. Then, we use those estimates to design the algorithm described in Section 5.2.
5.1 Confidence Sets and Maximal Matrix
Given the known SP order $\prec$ and user peaks $p(\cdot)$, we construct confidence sets for the expected reward matrix that contain the true parameters with high probability. Intuitively, these sets capture all matrices that are statistically plausible given the data observed so far, and naturally shrink as more data is collected. We begin by establishing notation for standard confidence bounds. Given a history $H_{t}$ at round $t$, we denote by $\bar{\theta}_{u,k}(t)$ and $n_{u,k}(t)$ the empirical mean reward and the number of pulls of the pair $(u,k)$ observed up to round $t$, respectively. The upper confidence bound is defined as $UCB_{u,k}(t)=\bar{\theta}_{u,k}(t)+\sqrt{\frac{2\ln T}{n_{u,k}(t)}}$, with the lower confidence bound $LCB_{u,k}(t)$ defined analogously using subtraction. We formally define the confidence set as follows.
Definition 4 (Confidence set).
The confidence set $\mathcal{C}^{t}=\mathcal{C}^{t}(\prec,p,H_{t})$ contains a matrix $P$ if and only if it is consistent with $\prec$ and $p(\cdot)$, and for any $u,k$ it holds that $LCB_{u,k}(t)\leq P_{u,k}\leq UCB_{u,k}(t)$.
The standard combinatorial UCB approach requires picking the optimistic matching in every time $t$, namely
This discrete optimization problem over a convex set may be tricky, as different matrices yield different optimal matchings. However, a structural property of $\mathcal{C}^{t}$ simplifies this optimization problem. As we demonstrate shortly in Lemma 7, $\mathcal{C}^{t}$ includes a unique maximal matrix $P^{t}$ in an element-wise sense; thus, to find the optimistic matching $\pi_{t}$, we should only find the optimal matching w.r.t. $P^{t}$.

### v2

@@ 6.1 Approximate Single-Peaked Matrices
We say a matrix $P$ is $\delta$-approximately single-peaked (or $\delta$-ASP) w.r.t. an order $\prec$ if for every $i,j,l\in[K]$ such that $i\prec j\prec l$ and user $u\in U$, it holds that $P_{u,j}\geq\min\{P_{u,i},P_{u,l}\}-\delta$.
@@
The $\delta$-ASP condition proves useful beyond the strict $\delta=0$ case. In fact, it follows directly from Definition˜4 that if $P$ is SP w.r.t. $\prec$ and $\tilde{P}$ satisfies $\lVert P-\tilde{P}\rVert_{\infty}\leq\delta$, then $\tilde{P}$ is $2\delta$-ASP w.r.t. $\prec$. This implies that a noisy estimate of an SP matrix remains ASP. Conversely, given any $\delta$-ASP matrix and its ASP order, we can construct a matrix that is (exact) SP with respect to the same order, with each entry differing from the original by at most $\delta$.
@@
We prove Lemma˜9 by construction, which can be implemented in $O(UK)$ time. Lemma˜9 enables us to work with SP matrices even when the estimate matrix is not SP, while not sacrificing the approximation quality. However, Lemma˜9 requires the ASP order $\prec$, which we do not have.
@@ 6.2 Extract-Order Procedure
We start with some intuition. Finding an order boils down to solving a set of contiguity constraints—subsets of arms that must be contiguous in any valid order. To illustrate, fix an SP matrix $P$ and suppose there exists a user $u$ and a partition of $[K]$ into two sets $A_{1}$ and $A_{2}$. Let $\beta=\min_{k\in A_{1}}P_{u,k}$ and $\gamma=\max_{k\in A_{2}}P_{u,k}$. If $\beta-\gamma>0$, then every arm in $A_{1}$ is preferred by user $u$ over every arm in $A_{2}$. In this case, any order $\prec$ that places an arm $j\in A_{2}$ between two arms $i,l\in A_{1}$ violates the 0-ASP condition (Definition˜4). Hence, according to Proposition˜8, $\prec$ cannot be an SP order of $P$. Thus, any SP order of $P$ places the arms in $A_{1}$ in a contiguous block. As a result, any order that satisfies all such contiguity constraints is an SP order of $P$.
@@
The Extract-Order algorithm, described in Algorithm˜2, formalizes this approach using a PQ tree [7], an efficient data structure that specializes in encoding and resolving contiguity constraints. PQ trees maintain sets of items that must appear consecutively in any valid order, and can be used both to determine whether a consistent order exists and to efficiently construct such an order. Extract-Order initializes a PQ tree (Line 3), and for each user $u$, it sorts the arms by preference values (Lines 4-5). Whenever there is a significant gap (exceeding $2\varepsilon$) between consecutive preference values, which exactly corresponds to the case described in ˜(4), it adds a contiguity constraint requiring all higher-valued arms to appear consecutively (Line 7). Finally, the algorithm returns an order on the columns of $\tilde{P}$ that satisfies all the imposed contiguity constraints, if such an order exists (Line 9). Otherwise, it fails and returns nothing (Line 11)
Notably, the constraints enforced by this procedure are not only consistent with the SP order of $P$, but also ensure bounded valley depth (in the sense of Definition˜4). Formally,
@@
Let $\tilde{P}$ be a matrix such that $\lVert\tilde{P}-P\rVert_{\infty}\leq\varepsilon$ for some SP matrix $P$. Extract-Order($\tilde{P},\varepsilon$) returns an order $\prec$ such that $\tilde{P}$ is $\left(2K\varepsilon\right)$-ASP w.r.t. $\prec$.
@@ 6.3 Explore-then-Commit Algorithm
4:Construct an SP matrix $\tilde{\Theta}$ from $(\bar{\Theta},\prec)$ via Lemma˜9 with $\delta=2K\sqrt{\nicefrac{{2\ln T}}{{N}}}$
@@
We now present the Explore-then-Match-and-Commit algorithm (EMC), described in Algorithm˜3, which combines the tools developed in the previous subsections to achieve sublinear regret. It operates in five phases. First, given an input parameter $N$, the algorithm explores by pulling each arm $k\in[K]$ for $N$ rounds and collects empirical means $\bar{\Theta}$. Second, it applies Extract-Order and extracts an order $\prec$ from $\bar{\Theta}$. The parameter $\varepsilon$ is chosen so that $\lVert\Theta-\bar{\Theta}\rVert_{\infty}\leq\varepsilon$ holds with high probability. In the third phase, it uses $\bar{\Theta}$ and the order $\prec$ to construct the matrix $\tilde{\Theta}$, which is SP w.r.t. $\prec$ and is element-wise close to $\bar{\Theta}$. Fourth, it executes SP-Matching on $\tilde{\Theta}$ to find an optimal matching $\tilde{\pi}$. And in the fifth and last phase, it exploits–it plays $\tilde{\pi}$ for the remaining TNKT-NK rounds. Importantly, the computational effort in Lines 35 is only $O(K^{2}U+K^{2}B)$, ensuring that the algorithm is computationally efficient. This is in sharp contrast to the exponential runtime we obtain for general instances in Section˜2.2.
The regret analysis for the commitment phase hinges on the approximation quality between the matrix $\tilde{\Theta}$ used for the matching and the true expected reward matrix $\Theta$. Since $\tilde{\Theta}$ is constructed to be close to the empirical estimates $\bar{\Theta}$, and $\bar{\Theta}$ concentrates around $\Theta$ with sufficient exploration, the value gap $V(\tilde{\pi};\tilde{\Theta})-V(\tilde{\pi};\Theta)$ is bounded by the estimation errors and the approximation errors from Lemma˜9 and Extract-Order. By optimizing the choice of exploration rounds $N$, we balance the exploration cost against the quality of the resulting approximation.
@@
6.4 Hardness of Optimistic Matching with Unknown Order
In the known SP structure regime (Section˜5), the MvM algorithm achieves $\tilde{O}(U\sqrt{TK})$ regret by solving the optimistic matching problem in each round. A natural question is whether a similar approach can yield improved regret guarantees when the SP structure is unknown. Recall that the optimistic approach requires solving Equation˜2, where $\mathcal{C}^{t}$ is now the confidence set of statistically plausible SP matrices. When the order and peaks are known, Lemma˜6 shows that $\mathcal{C}^{t}$ admits a unique element-wise maximal matrix, reducing the optimization to a single call to SP-Matching. However, when the order is unknown, $\mathcal{C}^{t}$ must include all SP matrices consistent with any valid order, and no such maximal element exists in general.
We note that if the optimistic matching could be solved efficiently in this regime, the same analysis as in Theorem˜7 would yield the same $\tilde{O}(U\sqrt{TK})$ regret guarantee. One might hope that a more sophisticated algorithm could achieve this. We show that this is unlikely by proving that even a simpler subproblem–finding the best matrix in $\mathcal{C}^{t}$ for a fixed matching–is NP-hard to approximate.
Theorem 12.
Consider the following optimization problem, termed Max-SP-WCS: given sets of users $U$ and arms $K$, a matching $\pi:U\to K$, and confidence intervals $\textnormal{LCB}_{u,k},\textnormal{UCB}_{u,k}]$ for each $(u,k)\in U\times K$, find $\max_{P\in\mathcal{C}}V(\pi;P)$, where $\mathcal{C}$ contains all SP matrices respecting the confidence intervals. Then, approximating Max-SP-WCS within a factor of $\frac{3}{4}+\delta$ is NP-hard for any $\delta>0$.
7 Extensions
In this section, we explore two extensions of our main results. First, we extend the MvM algorithm to handle instances with unknown structure in Section˜7.1. Second, we generalize the EMC algorithm to accommodate non-single-peaked instances in Section˜7.2. We defer the precise specifications and proofs of the algorithms to Appendix˜F, and provide overviews and proof sketches below.
7.1 Separated Instances
In this subsection, we improve our guarantees for instances that are statistically simpler to learn due to distinct preference values. Specifically, we assume separated instances, where we quantify separation by $\Delta_{\textnormal{global}}=\min_{u}\min_{i\neq j}\lvert\theta_{u,i}-\theta_{u,j}\rvert$. We introduce the Sep-MvM algorithm, which leverages this separation to improve upon the regret bound of EMC.
Proposition 13.
Sep-MvM yields a regret of at most
$\tilde{O}\left(U\min\left\{KT^{2/3},\sqrt{TK}+\nicefrac{{K}}{{\Delta_{\textnormal{global}}^{2}}}\right\}\right),$
for $\Delta_{\textnormal{global}}=\min_{u}\min_{i\neq j}\lvert\theta_{u,i}-\theta_{u,j}\rvert$, with a per-step computational complexity of $O(K^{2}U+K^{2}B)$.
Proof Sketch of Proposition˜13.
Sep-MvM employs a hybrid exploration-exploitation strategy. Initially, it explores all user-arm pairs in a round-robin fashion, maintaining confidence intervals $\mathcal{I}_{u,k}(t)=[\textnormal{LCB}_{u,k}(t),\textnormal{UCB}_{u,k}(t)]$ for every user-arm pair. The algorithm monitors for a separation event, defined as the point where, for each user, the confidence intervals of all arms become pairwise disjoint. A direct concentration analysis reveals that such separation must occur after at most $O(\nicefrac{{K\ln T}}{{\Delta_{\textnormal{global}}^{2}}})$ rounds.
The algorithm sets a timeout at $\tau=KT^{2/3}$ rounds. If separation is detected before $\tau$, the algorithm invokes Extract-Order to recover the SP order and subsequently switches to MvM using the estimated structural parameters. Conversely, if separation is not achieved by $\tau$, the algorithm infers that the instance is not separated and reverts to the standard EMC strategy. The final regret bound reflects the minimum of these two regimes. ∎
7.2 Beyond single-peaked instances
In this subsection, we extend EMC to work on non-single-peaked instances. Specifically, we design the NSP-EMC algorithm, which maintains the regret guarantees of EMC while incurring an additional penalty that scales with the distance from the nearest SP instance.
Proposition 14.
Fix a $\mathsf{CBR}$ instance $\mathcal{I}$ (possibly not SP), and let $\Theta$ be its expectation matrix. Then, NSP-EMC achieves a regret of at most $O\left(UKT^{2/3}(\ln T)^{1/3}+\gamma UKT\right)$, where $\gamma=\inf_{P\text{ is SP}}\lVert\Theta-P\rVert_{\infty}$, with a computational complexity of $O(UK^{2}\log(T)+K^{2}B)$.
Proof Sketch of Proposition˜14.
Fix an instance with a (possibly non-SP) expected reward matrix $\Theta$, and let $\gamma=\inf_{P\text{ is SP}}\lVert\Theta-P\rVert_{\infty}$. During the exploration phase of EMC, we collect empirical estimates $\bar{\Theta}$ satisfying $\lVert\bar{\Theta}-\Theta\rVert_{\infty}\leq\delta_{est}$ w.h.p., and therefore, $\bar{\Theta}$ lies within distance $\gamma+\delta_{est}$ of a valid SP matrix w.h.p. Consequently, invoking Extract-Order on $\bar{\Theta}$ with a parameter $\varepsilon\approx\gamma+\delta_{est}$ guarantees the recovery of a valid permutation, allowing us to project $\bar{\Theta}$ onto a $O(K(\gamma+\delta_{est}))$-nearby SP approximation $\tilde{\Theta}$. Using the exploration parameter and following the regret analysis from EMC, the algorithm achieves a regret of $\tilde{O}(UKT^{2/3})+O(UKT\gamma)$. Moreover, we do not need to know $\gamma$ in advance—NSP-EMC conducts a binary search over the possible parameters passed to Extract-Order, and finds the smallest one for which it does not fail. Then, we can relate this parameter implicitly to $\gamma$ and achieve the same regret bound. ∎
8 Discussion
We studied a budgeted matching problem under single-peaked preferences. While the general offline problem is NP-hard to approximate within a constant factor, imposing a single-peaked structure enables polynomial-time optimization. Nevertheless, we demonstrated that this structure does not trivialize the statistical aspect of the learning problem, which remains as challenging as the general preferences case. When the SP structure is unknown, we introduce an explore-then-commit approach that first extracts an approximate order, then commits to a near-optimal policy, yielding $\tilde{O}(UKT^{2/3})$ regret. For the more lenient case of known SP order and user peaks, we developed an optimistic algorithm based on maximal matrices that achieves $\tilde{O}(U\sqrt{TK})$ regret. We also conducted an experimental validation, which is deferred to Appendix H due to space constraints.
We identify several interesting directions for future work. First, comparing our algorithmic results with the lower bounds, we observe room for improvement. For the unknown SP structure regume, EMC achieves $\tilde{O}(UKT^{2/3})$ regret, while the $\Omega(U\sqrt{TK})$ lower bound is statistically tight (achievable by inefficient algorithms). It remains open whether the $\sqrt{T}$ rate can be achieved efficiently, or if our $T^{2/3}$ rate is tight for polynomial-time algorithms. Theorem˜12 provides evidence toward the latter: exponentially many SP structures may remain consistent with observations, and we show that even evaluating the optimistic value for a fixed matching is NP-hard to approximate. For known order and peaks, MvM achieves $\tilde{O}(U\sqrt{TK})$ regret (Theorem˜7), exceeding the lower bound by $\sqrt{\ln T}\cdot\max\{\sqrt{K},U\}$. Future work could resolve this gap.
Second, our results suggest that single-peaked preferences could potentially simplify other computationally challenging problems in online learning, suggesting an alternative route to a solution rather than $\alpha$-regret. Finally, extending our results to more complex preference structures [36, 32] presents an intriguing challenge.
Acknowledgments
This research was supported by the Israel Science Foundation (ISF; Grant No. 3079/24).
Appendix A General Preferences Analysis
In this appendix, we provide a comprehensive analysis of the computational complexity of achieving sublinear regret under general preferences. We first establish the fundamental hardness of the offline problem, then derive its implications for online learning, and finally discuss the $\alpha$-regret framework as a tractable alternative.
A.1 Hardness of the Offline Problem
We begin by proving that the offline budgeted matching problem is NP-hard to approximate within any factor better than (11/e)(1-1/e).
See 1
Proof of Theorem˜1.
We proceed via reduction from the Max $\ell$-Cover problem. In this problem, one is given a universe $\mathcal{U}$ of elements, a collection of subsets $\mathcal{S}=\{S_{1},\dots,S_{m}\}$, and an integer $\ell$. The objective is to select at most $\ell$ subsets whose union covers the maximum number of elements in $\mathcal{U}$. This problem is known to be NP-hard to approximate better than (11/e)(1-1/e) [18].
Given an instance of the Max $\ell$-Cover problem, we construct a $\mathsf{CBR}$ instance with $U=|\mathcal{U}|$ users (representing elements) and $K=m$ arms (representing subsets). We set unit costs $c_{k}=1$, budget $B=\ell$, and define preferences such that $\Theta_{u,k}=1$ if element $u\in S_{k}$ and 0 otherwise.
In this constructed instance, each matching $\pi$ induces a set of selected arms $\mathcal{K}=\{k:\exists u\in U\text{ s.t. }\pi(u)=k\}$. A user $u$ contributes $1$ to the total reward only if assigned to a selected arm $k$ where $u\in S_{k}$. Accordingly, given a fixed $\mathcal{K}$, the optimal matching allocates each user to a covering item if one exists; otherwise, the user yields zero reward. This implies that maximizing total reward is equivalent to maximizing coverage using at most $\ell$ items. It follows that any algorithm approximating the offline matching problem better than (11/e)(1-1/e) would violate the inapproximability result for Max $\ell$-Cover. ∎
A.2 Computational Barrier for Efficient Online Learning
The NP-hardness of approximating the offline problem beyond (11/e)(1-1/e) directly implies that efficient online learning with sublinear regret is impossible under standard complexity assumptions.
Corollary 15.
For any $U,K,B$ and $T=\mathrm{poly}(U,K,B)$, any poly-time online algorithm achieves regret $\Omega(T)$, unless $\mathrm{P}=\mathrm{NP}$.
Proof of Corollary˜15.
Restriction to hard instances.
By Theorem˜1, the inapproximability result holds already for instances produced by the reduction from Max-$\ell$-Cover [18]. These instances have $\Theta\in\{0,1\}^{U\times K}$ with unit costs, budget $B=\ell$, and at least one entry equal to $1$. Consequently, for such instances we always have $\max_{\pi\in\Pi(\Theta)}V(\pi;\Theta)\geq 1$. Since the hardness persists on this restricted subclass, it suffices to prove the corollary under the additional assumption $\max_{u,k}\theta_{u,k}=1$.
Construction of the online instance.
Simulating $\mathcal{A}$.
Run $\mathcal{A}$ for $T^{\prime}=U^{\alpha}D^{\beta}B^{\gamma}$ rounds on the expanded instance, obtaining matchings $\pi^{1},\ldots,\pi^{T^{\prime}}$, and return
$\pi^{\text{best}}=\arg\max_{t\in[T^{\prime}]}V(\pi^{t};\Theta^{\prime}).$
The regret guarantee yields
Approximation guarantee.
By construction of $D$, we ensured that $CT^{\prime\rho-1}\leq 0.1$. Hence
$V(\pi^{\text{best}};\Theta)\geq 0.9\,V(\pi^{\star};\Theta).$
Contradiction.
The algorithm above runs in $\text{poly}(U,K,B)$ time and produces a $0.9$-approximation to the offline optimum, contradicting the (11/e)(1-1/e)-hardness result from Theorem˜1 unless $\mathrm{P}=\mathrm{NP}$. ∎
A.3 Optimal Regret Without Computational Constraints
If we disregard computational constraints, optimal regret bounds can be attained by applying the standard CUCB algorithm [9], assuming access to an exact optimization oracle.
See 2
Proof of Corollary˜2.
The CUCB algorithm [9] maintains averages $\bar{\theta}_{u,k}(t)$ and counts $n_{u,k}(t)$ for each user-arm pair. In each round $t$, it computes UCB estimates $P_{u,k}(t)=\bar{\theta}_{u,k}(t)+\sqrt{\nicefrac{{2\ln T}}{{n_{u,k}(t)}}}$ (or $\infty$ if $n_{u,k}(t)=0$) and selects $\pi_{t}\in\operatorname*{arg\,max}_{\pi\in\Pi}V(\pi;P(t))$ using an optimization oracle. We note that we slightly changed the exact specification of the constants of CUCB, but it is of the same spirit.
By Hoeffding’s inequality and a union bound over all $(u,k)$ pairs and time steps, the clean event
$\mathcal{E}=\{\theta_{u,k}\in[\textnormal{LCB}_{u,k}(t),\textnormal{UCB}_{u,k}(t)]\text{ for all }u,k,t\}$
holds with probability at least $1-2UK/T^{3}$. We condition on $\mathcal{E}$; its complement contributes negligible regret.
Under $\mathcal{E}$, since $P(t)$ is element-wise maximal over all matrices respecting the confidence bounds, optimistic selection ensures that for any round $t$:
$\max_{\pi\in\Pi}V(\pi;\Theta)-V(\pi_{t};\Theta)$ $\leq V(\pi_{t};P(t))-V(\pi_{t};\Theta)$
$\leq\sum_{u\in[U]}2\sqrt{\frac{2\ln T}{n_{u,\pi_{t}(u)}(t)}}.$ (5)
The first inequality uses optimism: $V(\pi_{t};P(t))\geq V(\pi^{\star};P(t))\geq V(\pi^{\star};\Theta)$, where $\pi^{\star}$ is the optimal matching. The second inequality bounds the gap by the sum of confidence widths.
Summing Inequality 5 over all rounds and regrouping by arm pulls:
$R_{T}$ $\leq O\left(\sqrt{\ln T}\sum_{u\in[U]}\sum_{k\in[K]}\sum_{j=1}^{n_{u,k}(T)}\frac{1}{\sqrt{j}}\right)$
$=O\left(\sqrt{\ln T}\sum_{u\in[U]}\sum_{k\in[K]}\sqrt{n_{u,k}(T)}\right),$
where we used the standard bound $\sum_{j=1}^{n}1/\sqrt{j}\leq 2\sqrt{n}$. Applying Jensen’s inequality to the inner sum yields $\sum_{k}\sqrt{n_{u,k}(T)}\leq\sqrt{K\sum_{k}n_{u,k}(T)}=\sqrt{KT}$. Summing over all $U$ users gives $R_{T}=O(U\sqrt{TK\ln T})$.
The matching lower bound $\Omega(U\sqrt{TK})$ follows from Theorem˜3, as SP instances form a subclass of general instances. ∎
While this algorithm is statistically optimal, it assumes an optimization oracle that must solve an NP-hard problem at every time step, rendering it computationally impractical for general instances.
A.4 Efficient Approximation via Alpha-Regret
To bridge the gap between computational efficiency and performance guarantees, we consider the $\alpha$-regret framework, which compares against the best efficiently computable solution rather than the true optimum.
Definition 5 (Alpha Regret).